# Interactive-KBQA: Multi-Turn Interactions for Knowledge Base Question Answering with Large Language Models

### v1

@@ 4.2 Baselines
Fine-tuning (FT) on full data. We selected semantic parsing (SP)-based methods as baselines. DeCAF Yu et al. (2023) generates both logical forms and direct answers to get the final answers. For the KQA Pro, we opted for the BART-SPARQL Cao et al. (2022). For MetaQA, we choose Edge-aware Zhang et al. (2022b).
Prompting methods. We chose KB-BINDER Li et al. (2023a) which also leverage large language models (LLMs) in a few-shot setting. We have also compared the efficacy of Chain-of-Thought prompting (CoT prompt) Wei et al. (2022) alongside the technique of Self-Consistency (SC) Wang et al. (2023), further enriching our analysis. It’s important to note that StructGPTJiang et al. (2023b) and ToGSun et al. (2024) operate under the assumption that golden entities have been provided. Consequently, comparison of these methodologies is presented in Section 4.5.
FT on low-resource data. We fine-tuned open-source LLMs (open-LLMs) to directly generate SPARQL queries through supervised fine-tuning (SFT) manner, and reimplement a approach of prior SOTA in each dataset for comparison.
@@ 4.3 Evaluation Metrics
SP-based methods generate logical forms, which consequently produce answers in an unordered manner. These should be evaluated using the F1 score [^5] We use the average of the F1 scores across all instances.. However, we also report the Random Hits@1 (RHits@1) [^6] Randomly selecting an answer for each question 100 times and calculating the average Hits@1 Shu et al. (2022), and the Exact Match (EM) metricTalmor and Berant (2018) for reference. For KQA Pro, we report the accuracy, defined as the condition where the two sets match exactly.
@@ 4.4 Results
Table 2: Results on WebQSP and CWQ. Results tagged with $\dagger$ denote data from original prediction files, but evaluated on consistent test data. Results with $\sharp$ are reprinted from Li et al. (2023a).
@@
Table 2 and Table 3 present comprehensive comparisons. As indicated in the table, compared to methods that utilize full data, our approach is at a natural disadvantage on the WebQSP and KQA Pro datasets due to the significant difference in the magnitude of training data (~3K/~33K vs. 4-shot/2-shot, respectively). However, on CWQ and MetaQA, our approach overall outperforms baselines. Furthermore, in the cases of comparative and superlative question types, it achieves an improvement of 29.85% and 13.96%, respectively. This can be attributed to the fact that within the original dataset, these two types of questions each constitute merely 5% (see Appendix 11). Unlike previous methods, our approach is not constrained by data distribution, rendering it more robust.
@@ 4.5 Impact of Entity Linking (EL)
StructGPT $\dagger$ 44.26 - [^7] The output of StructGPT consists of a string that contains answers, making it unsuitable for evaluation with the RHits@1 metric. -
@@
The EL performance of our methods (with GPT-4 Turbo) is delineated in Table 4. Moreover, we introduce the Mention Cover Rate (MCR) to quantify the difficulty of EL, defined as the rate at which the golden entity names appear within the questions. As shown in our analysis, KQA Pro and MetaQA exhibit higher MCR. It is noted that some entity constrains are redundant in KQA Pro [^8] For instance, the query ”What is the connection between Steve Jordan (the one whose position is tight end) and Phoenix (the one that is the twinned administrative body of Chengdu)?” is equivalent to ”What is the connection between Steve Jordan and Phoenix?”.. Therefore, although the F1 score is 75.35%, this is not a bottleneck affecting the performance of KQA Pro. On the basis of these observations, we conducted experiments on the WebQSP and CWQ datasets under the conditions of given golden entities. The results, as presented in Table 5, reveal that our approach significantly outperforms the previous SOTA across both datasets.
@@ 4.6 Ablation Study
The performance of classifiers (cls) trained with our annotated dataset is shown in Table 6. Experiments were conducted to assess the effect of demonstrations in 0 and 4-shot scenarios on a randomly selected test subset of 100 entries. For CWQ, one exemplar per question type was used, while for KQA, only four types were sampled due to higher inference costs. In the zero-shot scenario, only the instruction text were used as prompt. As Table 7 shows, our method’s performance improves with increased question type coverage in prompts. For CWQ 4-shots, we observed a 2.5 point performance increase, but the cost rose by 37.86%, from \$0.5 to \$0.7. KQA Pro results suggest that accurate demonstrations improve performance and reduce costs. For a detailed discussion about the interaction rounds and costs, see Appendix A.5.
@@
As demonstrated in Table 8, GPT-4 significantly outperforms GPT-3.5101010We use gpt-3.5-turbo-1106. in terms of reasoning capabilities. We have also attempted to apply direct reasoning with Mistral 7B. It was observed that the untrained model significantly struggles with complex, multi-turn interactions. This finding underscores the substantial improvements that fine-tuning provides.
@@ 4.7 Error Analysis
Entity Linking Error refers to the failure to locate nodes using the SearchNodes tool. Predicate Search Error denotes the failure of the SearchGraphPatterns tool to return the necessary information. Reasoning Error means that, given the observations, the LLM fails to generate the appropriate SPARQL query. Format Compliance Error implies that the LLM does not use the tool in the required format. Hallucination includes generating elements that are inconsistent with the observations. Other Error encompasses errors that cannot be categorized under the above types.
More errors is presented in the form of case study in the Appendix A.7.
@@ 5 Conclusion
We present Interactive-KBQA, a KBQA framework employs an LLM as an agent to perform semantic parsing through multi-turn interactions with KBs. Initially, we developed a unified tool and interaction methodology tailored for various DB schemas. Subsequently, by categorizing complex questions and annotating a minimal set of exemplars, we employed a few-shot learning strategy that guides the LLM to incrementally generate SPARQL queries. Moreover, we introduced a low-resource dataset that demonstrates superior performance when fine-tuned with open-source LLMs.
@@ A.1 Additional Statistics of Datasets & Databases
Table 11: Distribution of (Dist.) question types in the original CWQ
Table 10 presents the statistical information for Freebase, Wikidata, and Movie KG. For Freebase, we utilize a script [^11] https://github.com/lanyunshi/Multi-hopComplexKBQA/blob/master/code/FreebaseTool/FilterEnglishTriplets.py to eliminate non-English triples. In the case of Wikidata, we refer to a dense subset processed by KQA Pro Cao et al. (2022). Meanwhile, Movie KG is derived by converting the knowledge base provided by MetaQA into RDF triples. Additionally, Table 11 outlines the distribution of question types in the original CWQ.
@@ A.2.1 Implementation Details
In the interactive process, our study invokes the OpenAI GPT4-Turbo gpt-4-1106-preview API to serve as a LLM agent. For each type of questions, we selected and annotated two cases with complete interactive processes as exemplars for in-context learning. During the inference stage, for the WebQSP and MetaQA datasets, all annotated demonstrations are employed as few-shot learning examples, specifically, 4-shot for WebQSP and 6-shot for MetaQA. For CWQ and KQA Pro, we train a bert-base-uncased  [^12] https://huggingface.co/bert-base-uncased model based on low-resource training data as a question type classifier, and then select two exemplars based on the predicted question type, namely cls + 2-shot for both CWQ and KQA Pro.
In the development of the SearchNodes tool, Elasticsearch [^13] https://github.com/elastic/elasticsearch is employed to extract all node surface names from the Freebase and MetaQA databases, and vector search techniques are implemented to perform queries on nodes within Wikidata. For the ranking algorithm of the SearchGraphPatterns tool, vector retrieval methods are likewise employed. All processes related to vectorization leverage the OpenAI text-embedding-ada-002 API. Moreover, for the functionality of the ExecuteSPARQL tool, Virtuoso [^14] https://github.com/openlink/virtuoso-opensource serves as the underlying graph query engine.
In the process of fine-tuning open LLMs, we utilize, Mistral-7B-v0.1 Jiang et al. (2023a) and Llama-2-13b Touvron et al. (2023). For training optimization, DeepSpeed Rasley et al. (2020) is employed, while inference tasks are accelerated using vLLM Kwon et al. (2023). Llama-2-13b is trained on four NVIDIA A100 80GB GPUs, the other experiments are conducted on two NVIDIA A100 80GB GPUs.
@@ A.2.2 Hyper-Parameter Setting
Table 12 presents the parameter configurations for invoking the OpenAI API and fine-tuned open-source LLMs. We employ DeepSpeed Rasley et al. (2020) to efficiently fine-tune open-source LLMs. The hyperparameter settings utilized for training are detailed in Table 13.
@@ A.5 Interaction Turns and Costs
In the process of interaction, the turn constitutes the fundamental unit, with each utterance made by any participant being recorded as a distinct turn. Table 18 provides a statistical analysis of the interaction turns. A case is deemed successful if the LLM agent explicitly generates Action:Done. Otherwise, the dialogue is concluded once the maximum turn number is reached. It is feasible to consider the turn count of human-annotated data as a lower bound. It is observed that, except for on CWQ, GPT-4 have approached this lower bound. Specifically, in CWQ, the overall success rate per turn for GPT-4 is 80.5/14.57, which, compared to WebQSP, shows a decrease of 14.66% and an increase of 39.96% in terms of success rate and turn count, respectively. This also delineates the difficulty level of the dataset from one perspective. In contrast, the KQA Pro dataset exhibits a stronger regularity, hence presenting a relatively lower level of difficulty. Furthermore, we discovered that on complicated question datasets, the average number of turns for fine-tuned LLMs is less than that for GPT-4, indicating that our annotated data effectively reduced the model’s trial-and-error process by providing valuable information. Additionally, we report the average cost per dialogue for reference  [^15] The price of gpt-4-1106-preview: Input \$0.01 / 1K tokens and Output \$0.03 / 1K tokens. Given that we set the return sequence parameter to 6 for inference and to 1 for human annotation, the cost comparison is not strictly equivalent. Complete results categorized by question type are available in the Table 20, Table 21 and Table 22.
@@
A.6 Confusion Matrix of Classifiers
@@
A.7 Case Study
In Figure 7, 8, 9, and 10, we present four cases. Each illustrates a error and delineates the correction process with human assistance.

### v2

@@ 4.2 Baselines
Fine-tuning (FT) on full data. We selected semantic parsing (SP)-based methods as baselines. DeCAF (Yu et al., 2023) generates both logical forms and direct answers to provide the final outcomes. For KQA Pro, we opted for the BART-SPARQL (Cao et al., 2022). For MetaQA, we choose Edge-aware (Zhang et al., 2022b).
Prompting methods. We chose KB-BINDER (Li et al., 2023a) which leverage large language models (LLMs) in a few-shot setting. Additionally, we compared the efficacy of Chain-of-Thought prompting (CoT prompt) (Wei et al., 2022) alongside the technique of Self-Consistency (SC) (Wang et al., 2023), further enriching our analysis. It’s important to note that StructGPT (Jiang et al., 2023b) and ToG (Sun et al., 2024) operate under the assumption that golden entities have been provided. Consequently, the comparison of these methodologies is presented in Section 4.5.
FT on low-resource data. We fine-tuned open-source LLMs (open-LLMs) to directly generate SPARQL queries through supervised fine-tuning (SFT) manner. We reimplemented a prior SOTA approach for each dataset to serve as a comparison.
@@ 4.3 Evaluation Metrics
SP-based methods generate logical forms, which consequently produce answers in an unordered manner. These should be evaluated using the F1 score [^6] We use the average of the F1 scores across all instances.. Additionally, we also report the Random Hits@1 (RHits@1) [^7] Randomly selecting an answer for each question 100 times and calculating the average Hits@1. (Shu et al., 2022), and the Exact Match (EM) metric (Talmor and Berant, 2018) for reference. For KQA Pro, we report the accuracy, defined as the condition where the two sets match exactly.
@@ 4.4 Results
Table 2: Results on WebQSP and CWQ. Results tagged with $\dagger$ denote data from original prediction files, but evaluated on consistent test data. Results with $\sharp$ are reprinted from (Li et al., 2023a).
@@
Table 2 and Table 3 present comprehensive comparisons. As indicated in the table, compared to methods that utilize full data, our approach is at a natural disadvantage on the WebQSP and KQA Pro datasets due to the significant difference in the magnitude of training data (~3K/~33K vs. 4-shot/2-shot, respectively). However, on CWQ and MetaQA datasets, our approach overall outperforms baselines. Furthermore, in the cases of comparative and superlative question types, our approach achieves improvements of 29.85% and 13.96%, respectively. This can be attributed to the fact that within the original dataset, these two types of questions each constitute merely 5% (see Table 11). Unlike previous methods, our approach is not constrained by data distribution, rendering it more robust. For conjunction questions in CWQ, CoT+SC and Decaf outperform our method due to redundant constraints inherent in these queries. Such questions require multiple entity constraints (e.g., "What country in the Mediterranean has Zonguldak Province?") yet often only one is needed to pinpoint the answer ("Zonguldak Province" leads to "Turkey", rendering "country in the Mediterranean" extraneous). Our approach, which focuses on identifying each predicate, complicates the response process.
@@ 4.5 Impact of Entity Linking (EL)
StructGPT $\dagger$ 44.26 - [^8] The output of StructGPT consists of a string that contains answers, making it unsuitable for evaluation with RHits@1. -
@@
The EL performance of our methods [^9] We find the entity surface name rather than mid. (with GPT-4 Turbo) is delineated in Table 4. The table reports a comparison of the results of our method with those of the widely recognized entity linking tool, ELQ (Li et al., 2020), on the WebQSP and CWQ datasets. Moreover, we introduce the Mention Cover Rate (MCR) to quantify the difficulty of EL. MCR is defined as the rate at which the golden entity names appear within the questions. As shown in our analysis, KQA Pro and MetaQA exhibit higher MCR. It is noted that some entity constrains are redundant in KQA Pro [^10] For instance, the answer of ”What is the connection between Steve Jordan (the one whose position is tight end) and Phoenix (the one that is the twinned administrative body of Chengdu)?” is equivalent to that of ”What is the connection between Steve Jordan and Phoenix?”.. Therefore, despite the F1 score being 75.35%, this is not a bottleneck affecting the performance of KQA Pro. Based on these observations, we conducted experiments on the WebQSP and CWQ datasets under the conditions of given golden entities. The results, as presented in Table 5, reveal that our approach significantly outperforms the previous SOTA across both datasets.
@@ 4.6 Ablation Study
This section conducts an ablation study to examine the influence of exemplars in two representative settings: CWQ, which includes four types and can barely cover all question types within the prompt text, and KQA Pro, which includes nine types and cannot cover all types due to high inference costs.
The performance of classifiers (cls) trained with our annotated dataset is shown in Table 6. Experiments were conducted in 0-shot and 4-shot scenarios on a randomly selected test subset of 100 entries. For CWQ, one exemplar per question type was used, while for KQA Pro, only four types were sampled. In the zero-shot scenario, only the instruction text was used as a prompt. As Table 7 shows, our method’s performance improves with increased question type coverage in prompts. For CWQ 4-shots, we observed a 2.5 point performance increase, but the cost rose by 37.86%, from \$0.5 to \$0.7. KQA Pro results suggest that accurate demonstrations improve performance and reduce costs. For a detailed discussion about the interaction rounds and costs, see Appendix A.5.
@@
As demonstrated in Table 8, GPT-4 significantly outperforms GPT-3.5121212We use gpt-3.5-turbo-1106. in terms of reasoning capabilities. We have also attempted to apply direct reasoning with Mistral 7B. It was observed that the untrained model significantly struggles with complex, multi-turn interactions. This finding underscores the substantial improvements that fine-tuning provides.
@@ 4.7 Error Analysis
  • Entity Linking Error refers to the failure to locate nodes using the SearchNodes tool. The primary cause of this error is the LLM’s failure to extract the correct entity mention from the question. For example, in the question "Most Anticipated Tour at the Young Hollywood Awards.", the entity should have been identified as "Young Hollywood Award for Most Anticipated Tour", but instead, the LLM mistakenly searched for "Young Hollywood Awards". Besides, In CWQ and WebQSP, the entities in the returned results do not have descriptions, leading the LLM to consider the results incorrect and subsequently re-initiate the search.
  • Predicate Search Error denotes the failure of the SearchGraphPatterns tool to return the necessary information. This error arises when vector search tools fail to return expected results, such as predicates indicating an organization’s headquarters. For instance, the query "locate in" is unable to match the predicate "organization.organization.headquarters".
  • Reasoning Error means that, given the observations, the LLM fails to generate the appropriate SPARQL query. This failure is largely due to an insufficient semantic understanding of the KG schema. Typically, the LLM often fails to understand the CVT structure (e.g. the direction of the predicate), resulting in incomplete or inaccurate SPARQL queries.
  • Format Compliance Error implies that the LLM does not use the tool in the required format. Examples include incorrectly constructed SPARQL queries with improperly formatted time and numerical values, and the introduction of unrecognized additional parameters to the tool.
  • Hallucination includes generating elements that are inconsistent with the observations.
  • Other Error encompasses errors that cannot be categorized under the above types.
More details are presented in the form of case study in the Appendix A.8.
@@ 5 Conclusion
Interactive-KBQA introduces a KBQA approach which utilizes an LLM as an agent for performing semantic parsing through multi-round interactions with a KB. Initially, we developed a unified tool and an interaction methodology tailored to various DB schemas. Subsequently, by categorizing complex questions and annotating a minimal set of exemplars, we employed a few-shot learning strategy that guides the LLM to incrementally generate SPARQL queries. Moreover, we introduced a low-resource dataset that demonstrates superior performance when fine-tuned with open-source LLMs.
@@
Acknowledgments
This work was supported by the National Key Research and Development Program of China under Grant No. 2023YFC3304404. We would also like to thank to Prof. Zhiyuan Liu and Dr. Yujia Qin from THUNLP group for their invaluable suggestions.
@@ A.1 Additional Statistics of Datasets & Databases
Table 11: Distribution (Dist.) of question types in the original CWQ.
Table 10 presents the statistical information for Freebase, Wikidata, and Movie KG. For Freebase, we utilize a script [^13] https://github.com/lanyunshi/Multi-hopComplexKBQA/blob/master/code/FreebaseTool/FilterEnglishTriplets.py to eliminate non-english triples. In the case of Wikidata, we refer to the subset processed by KQA Pro Cao et al. (2022). Meanwhile, Movie KG is derived by converting the knowledge base provided by MetaQA into RDF triples. Additionally, Table 11 outlines the distribution of question types in the original CWQ.
@@ A.2.1 Implementation Details
In the interactive process, our study invokes the OpenAI GPT4-Turbo (gpt-4-1106-preview) API to serve as a LLM agent. For each type of questions, we selected and annotated two cases with complete interactive processes as exemplars for in-context learning. During the inference stage, for the WebQSP and MetaQA datasets, all annotated demonstrations are employed as few-shot learning examples, specifically, 4-shot for WebQSP and 6-shot for MetaQA. For CWQ and KQA Pro, we train a bert-base-uncased  [^14] https://huggingface.co/bert-base-uncased model based on low-resource training data as a question type classifier (cls), and then select two exemplars based on the predicted question type, namely cls+2-shot for both CWQ and KQA Pro.
In the development of the SearchNodes tool, Elasticsearch [^15] https://github.com/elastic/elasticsearch is employed to extract all node surface names from the Freebase and MetaQA databases, and vector search techniques are implemented to perform queries on nodes within Wikidata. For the ranking algorithm of the SearchGraphPatterns tool, vector retrieval methods are similarly employed. All processes related to vectorization utilize the OpenAI text-embedding-ada-002 API to generate vectors and employ Chroma [^16] https://github.com/chroma-core/chroma for indexing and searching. Moreover, for the functionality of the ExecuteSPARQL tool, Virtuoso [^17] https://github.com/openlink/virtuoso-opensource serves as the underlying graph query engine.
In the process of fine-tuning open-source LLMs, we utilize, Mistral-7B-v0.1 Jiang et al. (2023a) and Llama-2-13b Touvron et al. (2023). For training optimization, DeepSpeed Rasley et al. (2020) is employed, while inference tasks are accelerated using vLLM Kwon et al. (2023). Llama-2-13b is trained on four NVIDIA A100 80GB GPUs, the other experiments are conducted on two NVIDIA A100 80GB GPUs.
@@ A.2.2 Hyper-Parameter Setting
Table 12 presents the parameter configurations for invoking the OpenAI API and fine-tuning open-source LLMs. We employ DeepSpeed Rasley et al. (2020) to efficiently fine-tune open-source LLMs. The hyperparameter settings utilized for training are detailed in Table 13.
@@ A.4 Additional Results of Exemplars Selection
The observation shows that the performance of our proposed approach on KQA Pro drops significantly in zero-shot compared to few-shot scenarios as observed in Table 17. This discrepancy arises primarily due to the intricate nature of the questions within the KQA Pro dataset, which involve multi-hop reasoning, constraints, and qualifiers. Constructed using SPARQL templates and phrased naturally by human annotators, these questions are complex. In zero-shot settings, without exemplars, even human annotators may struggle with semantic parsing due to the varied and unseen KG schema. Conversely, the few-shot approach leverages a small set of carefully annotated examples, effectively guiding the parsing process and accounting for the observed performance differences.
@@ A.5 Interaction Turns and Costs
In the process of interaction, the turn constitutes the fundamental unit, with each utterance made by any participant being recorded as a distinct turn. Table 18 provides a statistical analysis of the interaction turns. A case is deemed successful if the LLM agent explicitly generates Action:Done. Otherwise, the dialogue is concluded once the maximum turn number is reached. It is feasible to consider the turn count of human-annotated data as a lower bound. It is observed that, except for on CWQ, GPT-4 have approached this lower bound. Specifically, in CWQ, the overall success rate per turn for GPT-4 is 80.5/14.57, which, compared to WebQSP, shows a decrease of 14.66% and an increase of 39.96% in terms of success rate and turn count, respectively. This also delineates the difficulty level of the dataset from one perspective. In contrast, the KQA Pro dataset exhibits a stronger regularity, hence presenting a relatively lower level of difficulty. Furthermore, we discovered that on complicated question datasets, the average number of turns for fine-tuned LLMs is less than that for GPT-4, indicating that our annotated data effectively reduced the model’s trial-and-error process by providing valuable information. Additionally, we report the average cost per dialogue for reference  [^18] The price of gpt-4-1106-preview: Input \$0.01 / 1K tokens and Output \$0.03 / 1K tokens. Given that we set the return sequence parameter to 6 for inference and to 1 for human annotation, the cost comparison is not strictly equivalent. Complete results categorized by question type are available in the Table 20, Table 21, and Table 22.
@@
A.6 Impact of Interaction History Length
To investigate the impact of interaction history length on end-to-end performance, we analyzed the changes in F1 score as the number of dialogue turns increased, as presented in Figures 6. It was observed that as the number of dialogue turns increased, the model’s performance generally deteriorated.
Figure 6: F1 Scores across Different Turn Numbers.
A.7 Confusion Matrix of Classifiers
@@
A.8 Case Study
In Figure 8, 9, 10, and 11, we present four cases. Each illustrates a error and delineates the correction process with human assistance.